1. “Real” Analysis is a Degenerate Case of Discrete Analysis, Appeared in: ”New Progress in Difference Equations”(Proc. ICDEA 2001), edited by Bernd Aulbach, Saber Elaydi, and Gerry Ladas, and publisher by Taylor & Francis, London, 2004.

2. Towards a SymbolicComputational Philosophy (and Methodology!) for Mathematics, Proc. of the BuchbergerFest, (Peter Paule, ed.), Springer, to appear.

3. (With Stavros Garoufalidis and Thang TQ Le) The Quantum MacMahon Master Theorem, submitted.

4. Chomp, Recurrences, and Chaos, to Appear in J. Difference Equations and its Applications, J. Difference Equations and its Applications 10 (2004), 1281-1293. (special issue in honor of Saber Elaydi).

5. (With Xinyu Sun) On Fraenkel’s N-Heap Wythoff Conjecture, Annals of Combinatorics 8 (2004). 225-238.

6. (With Mohamud Mohammed) The Markov-WZ Method Elec J. Combinatorics 11(2004), R53. (14 pages).

7. Symbolic Moment Calculus I.: Foundations and Permutation Pattern Statistics, Appeared in Annals of Combinatorics 8 (2004), 369-378.

8. Dave Robbins’s Art of Guessing, Adv. Appl. Math. 34 (2005), 939-954. [special issue in memory of David Robbins].

9. (With Mohamud Mohammed) Sharp Upper Bounds for the Orders of the Recurrences Outputted by the Zeilberger and q-Zeilberger Algorithms, J. Symbolic Computation 39 (2005), 201-207.

10. Enumerative and Algebraic Combinatorics, to appear in: ‘Princeton Companion of Mathematics’, T.W. Gowers, ed.

11. (With Arthur Benjamin) Pythagorean Primes and Palindromic Continued Fractions, submitted.

12. (With Andrew V. Sills) Disturbing the Dyson Conjecture (in a GOOD Way), submitted.

13. DECONSTRUCTING the ZEILBERGER algorithm, to appear in J. of Difference Equations and Applications.

14. (With Mohamud Mohammed) Multi-Variable Zeilberger and Almkvist-Zeilberger Algorithms and the Sharpening of Wilf-Zeilberger Theory, submitted.